# Stochastic gradient descent-based inference for dynamic network models with attractors

### v2

@@ Abstract
In Coevolving Latent Space Networks with Attractors (CLSNA) models, nodes in a latent space represent social actors, and edges indicate their dynamic interactions. Attractors are added at the latent level to capture the notion of attractive and repulsive forces between nodes, borrowing from dynamical systems theory. However, CLSNA reliance on MCMC estimation makes scaling difficult, and the requirement for nodes to be present throughout the study period limit practical applications. We address these issues by (i) introducing a Stochastic gradient descent (SGD) parameter estimation method, (ii) developing a novel approach for uncertainty quantification using SGD, and (iii) extending the model to allow nodes to join and leave over time. Simulation results show that our extensions result in little loss of accuracy compared to MCMC, but can scale to much larger networks. We apply our approach to the longitudinal social networks of members of US Congress on the social media platform X. Accounting for node dynamics overcomes selection bias in the network and uncovers uniquely and increasingly repulsive forces within the Republican Party.
@@ 2.1 Model Definition
$\hbox{logit}(p_{t,ij})=\alpha-s(\bm{z}_{t,i},\bm{z}_{t,j})$ (2)
$Z_{t,i}\sim Normal(0,\tau^{2}I_{p})\enskip,$ (3)
and at time $t\geq 2$, if node $i$ is absent at time $t-1$,
$\hbox{logit}(p_{t,ij})=\alpha+\delta Y_{t-1,ij}-s(\bm{z}_{t,i},\bm{z}_{t,j})$ (4)
$Z_{t,i}\sim Normal(\mu_{t,i},\phi^{2}I_{p})$ (5)
@@
or else at time $t\geq 2$, if node $i$ is present at time $t-1$,
$\hbox{logit}(p_{t,ij})=\alpha-s(\bm{z}_{t,i},\bm{z}_{t,j})$ (7)
@@
Here $s(\cdot,\cdot)$ is a similarity function, and $A_{i}^{w}$ and $A_{i}^{b}$ are the two attractor functions for node $i$ in $Y_{t-1}$. Euclidean distance is used as the similarity function, $s(z_{t,i},z_{t,j})=||z_{t,i}-z_{t,j}||_{2}$, and the two attractors for node $i$ are defined as follows [^2] This application employs a two-group CLSNA model with attractors that mimic attractive and repulsive force. It is a specific version of the general CLSNA model class. Each node is assigned to one of two labeled groups, and movements are dictated by attractors shaped by neighboring nodes’ influence.,
$A_{i}^{w}(\bm{z}_{t-1},Y_{t-1})=\bar{\bm{z}}^{1}_{t-1,i}-\bm{z}_{t-1,i},\,\bar{\bm{z}}^{1}_{t-1,i}=\frac{1}{|S_{1}^{(i)}|}\sum_{j\in S_{1}^{(i)}}\bm{z}_{t-1,j}$ (10)
$A_{i}^{b}(\bm{z}_{t-1},Y_{t-1})=\bar{\bm{z}}^{2}_{t-1,i}-\bm{z}_{t-1,i},\,\bar{\bm{z}}^{2}_{t-1,i}=\frac{1}{|S_{2}^{(i)}|}\sum_{j\in S_{2}^{(i)}}\bm{z}_{t-1,j}\enskip,$ (11)
which are the discrepancies of $\bm{z}_{t-1,i}$ from two local averages at time $t-1$. The sets $S_{1}^{(i)}$ and $S_{2}^{(i)}$ are defined based on group membership and network connectivity as follows:
@@
Note that even though it is not explicitly mentioned in the notation, the sets $S_{1}^{(i)}$ and $S_{2}^{(i)}$ also depend on time because the connectivity sets are allowed to change with time. When $S_{1}^{(i)}$ or $S_{2}^{(i)}$ or both are empty, we set $A_{i}^{w}(\bm{z}_{t-1},Y_{t-1})=0$, or $A_{i}^{b}(\bm{z}_{t-1},Y_{t-1})=0$ or both to be zero respectively. This implies that if node $i$ is not connected to any members of a specific group, it receives no force from that group.
@@ 3.1 Point Estimate
      $\lambda$: step size for $Z_{1:T}$, $\lambda$: step size for $\theta$
@@
      M: maximum number of iterations, $\lambda$: step size for $Z_{1:T}$, $\lambda$: step size for $\theta$
@@
The resolution to this issue that we found to work well is to start with a CLSNA model with latent variables living in a higher dimensional space from the one we intend on having, that is, we aim to fit a model with $Z\in\mathbb{R}^{p}$ but we intentionally choose $Z\in\mathbb{R}^{q}$ with $q>p$ Then, we use the first $p$ principal components of the vector of the latent position variables for $Z_{1:T}$. This method was the most effective of several alternatives attempted to remedy local modes. Intuitively, one can think of this approach as creating a path between modes thus allowing SGD to travel between modes in the higher dimension. In both the simulated and real data examples, we initially set $p$ to 2 and chose $q$ as 3. We begin by fitting a model with $Z$ in $\mathbb{R}^{3}$, followed by applying principal component analysis (PCA) to the resulting latent positions, thereby compressing $Z$ into $\mathbb{R}^{2}$ while maximizing information preservation. This reduced form is then used as the initial value for fitting a model with $Z$ in $\mathbb{R}^{2}$. This approach, albeit doubling the cost of fitting the model and performing PCA, provides more accurate and faster convergence towards the true values, as it effectively navigates through potential local optima that could hinder convergence in a single-stage fitting process.
@@ 3.2 Variance Estimation
Under an MCMC framework, both point estimates for $\theta$ and corresponding uncertainties can be estimated using samples drawn from the posterior. However, our proposed SGD approach above only gives a point estimate for $\theta$, without quantifying uncertainty.
We propose a novel approach to estimate the posterior variance for parameters of interest. It is a method for approximate Bayesian inference based on Laplace’s method. Laplace’s approximation, when applied in scenarios like state-space models and Gaussian Markov Random Fields, leverages the sparsity of the precision matrix as detailed by Rue et al. (2009). However, in more general contexts, such as CLSNA, this assumption of sparsity is not always valid. In response to this, our proposed approach builds upon Laplace’s foundational principles but adapts to the challenges of non-sparse structures. Specifically, it transforms the variance estimation problem into an optimization problem that can also be solved with SGD.
In this section, we present the novel variance estimation algorithm that we propose.
3.2.1 Some Useful Results of Multivariate Normal Distribution
More specifically, the covariance of the conditional distribution $\text{Cov}(x_{1}|x_{2})$ is a constant as a function of $x_{2}$, indicating that the shape of the conditional distribution is independent of the specific value of $x_{2}$. Additionally, writing $p(x_{1}|x_{2})$ as the p.d.f of the conditional distribution of $x_{1}|x_{2}$, the mean $\mu_{1|2}=\mu_{1}+\Sigma_{12}\Sigma_{22}^{-1}(x_{2}-\mu_{2})$ is the expectation, median, and mode of the distribution. The maximum of the p.d.f. function $p(x_{1}=\mu_{1|2}|x_{2})$ is a constant as a function of $x_{2}$.
Theorem 3.1.
The proof of the theorem is provided in the supplementary file. In particular, the marginal distribution is proportional to the joint distribution evaluated on the curve of the conditional mean of $x_{1}$ given $x_{2}$. Theorem 7.1 provides an expression for the shape of the marginal distribution from the joint distribution, when the joint distribution is a multivariate normal distribution.
Next, we proceed with an important result demonstrating that for a normal distribution the ideas of Theorem 7.1 yield a formula for the variance of $x_{2}$ that we can then be turned into a practical SGD-based algorithm for uncertainty quantification, see Subsection 3.2.2. In particular, recall that in a univariate normal distribution, two key parameters - the mean and variance - describe the distribution fully. By considering two distinct points from the distribution, we can deduce these parameters. In Figure 2 we aim to visualize this idea, which then leads us to Corollary 7.1.1 containing the expression for the variance.
$x$$y$$p(x_{2})$“The curve corresponding to $(\mu_{1|2},x_{2})$$p(x_{1},x_{2})$$(\mu_{2},p(\mu_{1},\mu_{2}))$$(\mu_{2}+\eta,p(\mu_{1|2},\mu_{2}+\eta))$
Figure 2: Variance Estimation in Bivariate Normal Distribution
Corollary 3.1.1.
[…]
The proof of the corollary is provided in the supplementary file.
Corollary 3.1.2.
$\Sigma_{12}=\frac{\Sigma_{22}}{(\tilde{x}_{2}-\mu_{2})}(\mu_{1|2}-\mu_{1}),$
The methodologies established for a generic normal distribution $p(x_{1},x_{2})$ are now adapted to estimate variance in the posterior distribution $\pi(\cdot|Y)$, effectively applying the same theoretical concepts to the Bayesian analysis under a Laplace approximation.
@@ 3.2.2 The Variance Estimation Algorithm
Corollary 7.1.1 and Corollary 3.1.2 are the primary components in our proposed SGD-based method for variance estimation of a single parameter, as detailed in Algorithm 3. The method involves two stages. Initially, we compute a point estimate using an SGD parameter estimation method. In the second stage, with this point estimate, we will redo the optimization with the constraint of fixing the parameter to be variance-estimated, adjusting it by a certain $\eta$ perturbation size. Thanks to Corollary 7.1.1 and Corollary 3.1.2, we can then estimate the variance/covariance of this parameter by observing the changes in the posterior function and the changes in the other parameters resulting from the perturbation.
Input: Network time-series $Y_{1:T}$. The mode of $\log\pi(Z,\theta|Y)$, $\theta^{*}$ denoting the mode for all hyperparameters $(\alpha,\delta,\gamma^{w},\gamma^{b})$, $Z_{1:T}^{*}$ for latent positions.
      M: maximum number of iterations, $\lambda$: step size for $Z_{1:T}$, $\lambda$: step size for $\theta$, $\eta_{\alpha}$: perturbation size for $\alpha$       Output: $\widehat{\text{Var}(\alpha|Y)}$(the posterior variance of $\alpha$) (where $\alpha$ is the intercept term in Eq. 2)
1:Set initial values $\theta^{(0)}$, $Z_{1:T}^{(0)}$ to be $\theta^{*}$, $Z_{1:T}^{*}$.
@@
3:     At the onset of each iteration, the parameter of interest is anchored at a position that’s a fixed perturbation, denoted as $\eta_{\alpha}$, from its modal value. Specifically, we set $\alpha^{(k)}=\alpha^{*}+\eta_{\alpha}$. This action positions the tuple $(\alpha^{(k)},\delta^{(k)},\gamma^{w,(k)},\gamma^{b,(k)})$ at $(\alpha^{*}+\eta_{\alpha},\delta^{(k)},\gamma^{w,(k)},\gamma^{b,(k)})$.
4:     Take a SGD step to optimize the posterior distribution: $(\theta^{(k+1)},Z_{1:T}^{(k+1)})$ = SGDstep($\theta^{(k)},Z_{1:T}^{(k)}$)
5:     Stop if stopping criteria is met and let $k$ be the corresponding iteration.
@@
7:
$\widehat{\text{Var}(\alpha|Y)}=\frac{1}{2}\frac{\eta_{\alpha}^{2}}{\pi(\theta^{(0)},Z^{(0)}_{1:T}|Y_{1:T})-\pi(\theta^{(k+1)},Z^{(k+1)}_{1:T}|Y_{1:T})}$
8:
$\widehat{\text{Cov}(\alpha,\theta_{i}|Y)}=\frac{\widehat{\text{Var}(\alpha|Y)}}{\eta_{\alpha}}(\theta_{i}^{(k+1)}-\theta_{i}^{(0)})$
9:
$\widehat{\text{Cov}(\alpha,Z_{1:T}|Y)}=\frac{\widehat{\text{Var}(\alpha|Y)}}{\eta_{\alpha}}(Z_{1:T}^{(k+1)}-Z_{1:T}^{(0)})$
$\text{Cov}(\alpha,Z_{1:T}|Y)$ denote the covariance vector between the 1-dimensional random variable $\alpha$ and the vector random variable $Z_{1:T}$.
Algorithm 3 details an SGD-based approach for estimating the variance of a single parameter, in this instance for $\alpha$. To extend this estimation to all parameters of interest, we repeatedly apply Algorithm 3 individually to each parameter. Since we are able to recover the covariance matrix, we are also able to estimate variance of functions of parameters. As an illustration, let us look for example into $\text{Var}(\gamma^{w}-\gamma^{b})$. This estimation is achieved through a sequential application of Algorithm 3 to each component parameter. Initially, the algorithm is applied to $\gamma^{w}$, yielding estimates $\widehat{\text{Var}(\gamma^{w}|Y)}$ and $\widehat{\text{Cov}(\gamma^{w},\gamma^{b}|Y)}$. Following this, a similar application to $\gamma^{b}$ provides $\widehat{\text{Var}(\gamma^{b}|Y)}$ and an alternative estimate of $\widehat{\text{Cov}(\gamma^{w},\gamma^{b}|Y)}$. The variance of the composite parameter $\gamma^{w}-\gamma^{b}$ is then approximated using the formula $\widehat{\text{Var}(\gamma^{w}|Y)}+\widehat{\text{Var}(\gamma^{b}|Y)}-2\widehat{\text{Cov}(\gamma^{w},\gamma^{b}|Y)}$. Given that the function is not exactly quadratic, discrepancies between the two covariance estimates may arise. To mitigate this, one can proceed with averaging these estimates for a more robust estimation.

### v3

@@ Abstract
In Coevolving Latent Space Networks with Attractors (CLSNA) models, nodes in a latent space represent social actors, and edges indicate their dynamic interactions. Attractors are added at the latent level to capture the notion of attractive and repulsive forces between nodes, borrowing from dynamical systems theory. However, CLSNA reliance on MCMC estimation makes scaling difficult, and the requirement for nodes to be present throughout the study period limit practical applications. We address these issues by (i) introducing a Stochastic gradient descent (SGD) parameter estimation method, (ii) developing a novel approach for uncertainty quantification using SGD, and (iii) extending the model to allow nodes to join and leave over time. Simulation results show that our extensions result in little loss of accuracy compared to MCMC, but can scale to much larger networks. We apply our approach to the longitudinal social networks of members of US Congress on the social media platform X. Accounting for node dynamics overcomes selection bias in the network and uncovers uniquely and increasingly repulsive forces within the Republican Party. Supplemental materials for the article are available online.
@@ 2.1 Model Definition
$\hbox{logit}(p_{t,ij})=\alpha-s(z_{t,i},z_{t,j})$ (2)
$Z_{t,i}\sim Normal(0,\tau^{2}I_{p})\enskip.$ (3)
If, at time $t\geq 2$, if node $i$ was absent from the network at time $t-1$,
$\hbox{logit}(p_{t,ij})=\alpha+\delta Y_{t-1,ij}-s(z_{t,i},z_{t,j})$ (4)
$Z_{t,i}\sim Normal(\beta_{t,i}Z_{t_{0},i}+(1-\beta_{t,i})\mu_{t,i},\phi^{2}I_{p})$ (5)
@@
where $\beta_{t,i}\in[0,1]$ and $Z_{t_{0},i}$ denotes the potential position of the specific node $i$ at the most recent past time $t_{0}<t-1$ that it was present at the network. If node i has never been in the network before time $t$, then we simply set $\beta_{t,i}=0$. See also Remark 2.1.
If at time $t\geq 2$, node $i$ was present at the network at time $t-1$,
$\hbox{logit}(p_{t,ij})=\alpha-s(z_{t,i},z_{t,j})$ (7)
@@
Here $s(\cdot,\cdot)$ is the Euclidean distance, $s(z_{t,i},z_{t,j})=||z_{t,i}-z_{t,j}||_{2}$. $A_{i}^{w}$ and $A_{i}^{b}$ are the within group $\pi(i)$ (the group node $i$ belongs to) and between groups attractor functions for node $i$ in $Y_{t-1}$ respectively. In addition, $\gamma^{\pi(i)}_{w}$ and $\gamma_{b}$ are the attraction parameters within group $\pi(i)$ and between groups attraction parameters respectively. The two attractors for node $i$ are defined as follows [^2] This application employs a two-group CLSNA model with attractors that mimic attractive and repulsive force. It is a specific version of the general CLSNA model class. Each node is assigned to one of two labeled groups, and movements are dictated by attractors shaped by neighboring nodes’ influence.,
$A_{i}^{w}(z_{t-1},Y_{t-1})=\bar{z}^{1}_{t-1,i}-z_{t-1,i},\,\bar{z}^{1}_{t-1,i}=\frac{1}{|S_{1}^{(i)}|}\sum_{j\in S_{1}^{(i)}}z_{t-1,j}$ (10)
$A_{i}^{b}(z_{t-1},Y_{t-1})=\bar{z}^{2}_{t-1,i}-z_{t-1,i},\,\bar{z}^{2}_{t-1,i}=\frac{1}{|S_{2}^{(i)}|}\sum_{j\in S_{2}^{(i)}}z_{t-1,j}\enskip,$ (11)
which are the discrepancies of $z_{t-1,i}$ from two local averages at time $t-1$. The sets $S_{1}^{(i)}$ and $S_{2}^{(i)}$ are defined based on group membership and network connectivity as follows:
@@
Note that even though it is not explicitly mentioned in the notation, the sets $S_{1}^{(i)}$ and $S_{2}^{(i)}$ also depend on time because the connectivity sets are allowed to change with time. When $S_{1}^{(i)}$ or $S_{2}^{(i)}$ or both are empty, we set $A_{i}^{w}(z_{t-1},Y_{t-1})=0$, or $A_{i}^{b}(z_{t-1},Y_{t-1})=0$ or both to be zero respectively. This implies that if node $i$ is not connected to any members of a specific group, it receives no force from that group.
Remark 2.1.
We remark that (5) allows to cover the case where a given node $i$ leaves the network at a given time $t_{0}$ and re-enters again at a future time $t>t_{0}+1$. In that case, (5) shows that the mean of a new position can be defined to be a convex combination of the last known position and the average of the position of the rest of the nodes within the group. If no further information exists, taking $\beta_{t,i}=\frac{1}{2}$ is a reasonable choice. Of course, if node $i$ enters the network for the first time at time $t$, then the intuitive choice is to take $\beta_{t,i}=0$.
@@
We also present a graphical representation of the proposed model in Figure 2. The graphical representation is adapted from Zhu et al. (2023). This diagram illustrates the key dependencies in the model. The arrow from $Y_{t-1,ij}$ to $Y_{t,ij}$ represents the edge persistence effect described by Eq. (4), indicating that the presence of an edge at time $t-1$ influences the likelihood of an edge at time $t$. The arrow from $Y_{t-1,ij}$ to $Z_{t}$ reflects the influence of prior network connectivity on the evolution of latent positions, as captured by Eq. (8). The generation of edges based on latent position similarity, as described in Eq. (1), is depicted by the arrows from $Z_{t}$ to $Y_{t,ij}$. Finally, the arrow from $Z_{t-1}$ to $Z_{t}$ illustrates the influence of the previous position, i.e., at time $t-1$ on the position at time $t$.
[…]
Figure 2: Graphical model representation of the CLSNA model. All key dependencies are shown, including edge generation, edge persistence and temporal evolution
@@ 3.1 Point Estimate
      $\lambda$: step size for $Z_{1:T}$ and for $\theta$
@@
      M: maximum number of iterations, $\lambda$: step size for $Z_{1:T}$ and for $\theta$
@@
The resolution to this issue that we found to work well is to start with a CLSNA model with latent variables living in a higher dimensional space from the one we intend on having; that is, we aim to fit a model with $Z\in\mathbb{R}^{p}$ but we intentionally choose $Z\in\mathbb{R}^{q}$ with $q>p$ Then, we use the first $p$ principal components of the vector of the latent position variables for $Z_{1:T}$. This method was the most effective of several alternatives attempted to remedy local modes. Intuitively, one can think of this approach as creating a path between modes thus allowing SGD to travel between modes in the higher dimension. In both the simulated and real data examples, we initially set $p$ to 2 and chose $q$ as 3. We begin by fitting a model with $Z$ in $\mathbb{R}^{3}$, followed by applying principal component analysis (PCA) to the resulting latent positions, thereby compressing $Z$ into $\mathbb{R}^{2}$ while maximizing information preservation. This reduced form is then used as the initial value for fitting a model with $Z$ in $\mathbb{R}^{2}$. This approach, albeit doubling the cost of fitting the model and performing PCA, provides more accurate and faster convergence towards the true values, as it effectively navigates through potential local optima that could hinder convergence in a single-stage fitting process.
@@ 3.2 Variance Estimation
The first stage with SGD above only gives point estimates, without quantifying uncertainty. We propose a novel approach to estimate the posterior variance for parameters of interest. It is a method for approximate Bayesian inference based on Laplace’s method. Laplace’s approximation, when applied in scenarios like state-space models and Gaussian Markov Random Fields, leverages the sparsity of the precision matrix as detailed by Rue et al. (2009). However, in more general contexts, such as CLSNA, this assumption of sparsity is not always valid. In response to this, our proposed approach builds upon Laplace’s foundational principles but adapts to the challenges of non-sparse structures. Specifically, it transforms the variance estimation problem into an optimization problem that can also be solved with SGD.
In this section, we propose a novel variance estimation algorithm. Our variance estimation method exploits the properties of the conditional distribution of a multivariate Gaussian distribution. Let $x$ follow a multivariate normal distribution $x\sim\mathcal{N}(\mu,\Sigma)$, where the block-wise mean and covariance matrices are defined as $\mu=\left[\begin{smallmatrix}\mu_{1}\\ \mu_{2}\end{smallmatrix}\right]$ and $\Sigma=\left[\begin{smallmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{smallmatrix}\right]$. The conditional distribution of a subset vector $x_{1}$, given its complement vector $x_{2}$, is also a multivariate normal distribution $x_{1}|x_{2}\sim\mathcal{N}(\mu_{1|2},\Sigma_{1|2})$. We will use $\mu_{1|2}(x_{2})$ to denote the conditional mean of $x_{1}$ given $x_{2}$. With this notation, the variance and covariance can be respectively calculated as:
$\text{Var}(x_{2})$ $=\frac{1}{2}\frac{(\mu_{2}-\tilde{x}_{2})^{2}}{\log p(x_{1}=\mu_{1},x_{2}=\mu_{2})-\log p(x_{1}=\mu_{1|2}(\tilde{x}_{2}),x_{2}=\tilde{x}_{2})},$ (15)
$\Sigma_{12}$ $=\frac{\Sigma_{22}}{(\tilde{x}_{2}-\mu_{2})}(\mu_{1|2}(\tilde{x}_{2})-\mu_{1}).$ (16)
A more detailed discussion and proof of these properties can be found in the supplementary file. The results above for a generic normal distribution $\mathcal{N}(\mu,\Sigma)$ are now adapted to estimate variance in the posterior distribution $\pi(\cdot|Y)$, effectively applying the same theoretical concepts to the Bayesian analysis under a Laplace approximation.
@@ 3.2.1 The Variance Estimation Algorithm
Eq.17 and Eq.16 are the primary components in the second stage of our proposed SGD-based method for variance estimation of a single parameter, as detailed in Algorithm 3. The first stage detailed in section 3.1 computes a point estimate using SGD. In the second stage, with this point estimate, we redo the optimization with the constraint of fixing the parameter to be variance-estimated, adjusting it by a certain $\eta$ perturbation size. Thanks to Eq.17 and Eq.16, we can then estimate the variance/covariance of this parameter by observing the changes in the posterior function and the changes in the other parameters resulting from the perturbation.
Algorithm 3 details the second stage of the SGD-based approach for estimating the variance of a single parameter (in here we choose $\alpha$).
Input: Network time-series $Y_{1:T}$; mode of $\log\pi(Z,\theta|Y)$, $\theta^{*}$ (for all hyperparameters $\alpha,\delta,\gamma^{w},\gamma^{b}$), $Z_{1:T}^{*}$ (latent positions); $M$: max iterations; $\lambda$: step size for $Z_{1:T}$ and for $\theta$; $\eta_{\alpha}$: perturbation size for $\alpha$.
Output: $\widehat{\text{Var}(\alpha|Y)}$, $\widehat{\text{Cov}(\alpha,\theta_{i}|Y)}$, $\widehat{\text{Cov}(\alpha,Z_{1:T}|Y)}$
1:Set initial values $\theta^{(0)}$, $Z_{1:T}^{(0)}$ to $\theta^{*}$, $Z_{1:T}^{*}$.
@@
3:     Set $\alpha^{(k)}=\alpha^{*}+\eta_{\alpha}$.
4:     $(\theta^{(k+1)},Z_{1:T}^{(k+1)})=\textsc{SGDstep}(\theta^{(k)},Z_{1:T}^{(k)})$.
5:     Stop if criteria are met; let $k$ be the corresponding iteration.
@@
7:$\widehat{\text{Var}(\alpha|Y)}=\frac{1}{2}\frac{\eta_{\alpha}^{2}}{\log\pi(\theta^{(0)},Z^{(0)}_{1:T}|Y_{1:T})-\log\pi(\theta^{(k+1)},Z^{(k+1)}_{1:T}|Y_{1:T})}$
8:$\widehat{\text{Cov}(\alpha,\theta_{i}|Y)}=\frac{\widehat{\text{Var}(\alpha|Y)}}{\eta_{\alpha}}(\theta_{i}^{(k+1)}-\theta_{i}^{(0)})$
9:$\widehat{\text{Cov}(\alpha,Z_{1:T}|Y)}=\frac{\widehat{\text{Var}(\alpha|Y)}}{\eta_{\alpha}}(Z_{1:T}^{(k+1)}-Z_{1:T}^{(0)})$
In Figure 3 we aim to visualize this approach for a a Bivariate Distribution.
$x$$y$$\widehat{\pi(\alpha|Y)}$“The curve corresponding to $(X^{*}_{\alpha},\alpha)$$\pi(X,\alpha|Y)$$(\alpha,\pi(X^{*},\alpha|Y))$$(\alpha+\eta,\pi(X^{*}_{\alpha=\alpha+\eta},\alpha+\eta|Y))$
Figure 3: Illustration of the Variance Estimation Method in a Bivariate Distribution
To extend this estimation to all parameters of interest, we repeatedly apply Algorithm 3 individually to each parameter. Since we are able to recover the covariance matrix, we are also able to estimate variance of functions of parameters. As an illustration, let us look for example into $\text{Var}(\gamma^{w}-\gamma^{b})$. This estimation is achieved through a sequential application of Algorithm 3 to each component parameter. Initially, the algorithm is applied to $\gamma^{w}$, yielding estimates $\widehat{\text{Var}(\gamma^{w}|Y)}$ and $\widehat{\text{Cov}(\gamma^{w},\gamma^{b}|Y)}$. Following this, a similar application to $\gamma^{b}$ provides $\widehat{\text{Var}(\gamma^{b}|Y)}$ and an alternative estimate of $\widehat{\text{Cov}(\gamma^{w},\gamma^{b}|Y)}$. The variance of the composite parameter $\gamma^{w}-\gamma^{b}$ is then approximated using the formula $\widehat{\text{Var}(\gamma^{w}|Y)}+\widehat{\text{Var}(\gamma^{b}|Y)}-2\widehat{\text{Cov}(\gamma^{w},\gamma^{b}|Y)}$. Given that the function is not exactly quadratic, discrepancies between the two covariance estimates may arise. To mitigate this, one can proceed with averaging these estimates for a more robust estimation.